# Automatic gradient descent with generalized Newton’s method

### v1

@@ 3.4 Error analysis for optimal learning rate
We now analyze the estimation error of $\eta^{*}$ in (3.3), with batch size $B$ and mini-batch loss $\bar{L}(\bm{w}_{t})$. The estimation error consists of the sub-sampling error by using the mini-batch, and the precision error from fitting (2.3).
@@ 4 Efficiency and Scalability Analysis
In this section, we show that GeN optimizers are as computationally efficient and scalable (in terms of the utility and the system design) as their base optimizers. We train CIFAR10 [34] on ResNet 18, 34, 50, 152 [24] and ViT tiny, small, base and large [17]. For fine tuning, we use the pretrained models from the PyTorch Image Models framework [70].
@@
To set the stage, GeN optimizers clearly have the same peak memory cost as the base optimizers, as the latter also uses the forward pass. Hence it suffices to count the additional time complexity introduced by GeN. In the default setting, we can write
@@
where ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}F}$ is the time complexity of the forward pass, in terms of float-point operations, ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}B}$ is that of the back-propagation, and $C$ is other costs including the communication and the data loading, which are minimal on single GPU. Given that in full-parameter training, ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}B}\approx 2{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}F}$ [^5] Forward pass takes 1 unit of time. Back-propagation consists of two sub-processes – output gradient and parameter gradient, each taking 1 unit of time. See the complexity analysis in Table 1 of [10]., GeN optimizers are roughly 60% as fast as the base optimizers.
@@ 4.1 Lazy learning rate update
One simple trick to accelerate GeN is to update the learning rate every $\Phi$ iterations for $\Phi>1$, so the additional computation is amortized. For instance, GeN achieves $86\%$ relative speed at $\Phi=4$ and $>92\%$ relative speed at $\Phi\geq 8$. Empirically, applying $\Phi$ has insignificant effect on the convergence.
@@
In the PET setting, most parameters are frozen whose gradients are not computed. Hence ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}B}\approx{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}F}$ and the relative speed of GeN becomes $\frac{1}{1+1/\Phi}$, e.g. 50% at $\Phi=1$ and 89% at $\Phi=8$. That is, GeN has slightly lower relative speed than the full-parameter training, but enjoys a higher absolute speed.
@@ 4.2 Communication in distributed learning
For large-scale optimization, the distributed learning with multiple devices is necessary but challenging, in that (1) the communication orchestra is complicated, and (2) the communication cost is significant ($C\gg 0$). These two challenges determine if and how well an optimizer scales under a distributed solution. For data-parallel solutions such as Distributed Data Parallel and ZERO1/2 [57], the communication orchestra is relatively simple so that many Hessian-related optimizers are executable. For model-parallel and pipeline-parallel solutions such as ZERO3 and FSDP, each forward pass requires the communication of model parameters, and GeN indeed adds a significant amount of communication volume. Nevertheless, applying the lazy learning rate can essentially reduce the communication overhead to a negligible level.
@@ 6 Experiments on real data
In this section, we test GeN on real datasets across image classification, text classification, natural language generation, object detection and instance segmentation. We additionally experiment on image generation and recommendation system in Appendix B, where all training details can be found. Our experiments cover end-to-end training as well as fine-tuning, and include PET methods such as LoRA [28] and BitFit [72].
@@ 6.1 Image classification
reference [34] [7] [27] [48] [74] [1]
@@ 6.2 Natural Language Processing
For NLG experiments, we train GPT2 [55] model with LoRA on the E2E dataset [49]. We measure the quality of the generated texts by the perplexity (the exponential of loss), BLEU [50], ROUGE-L [35], METEOR [5], NIST [15], and CIDEr [67]. We use [28] as the baseline, which trains for 5 epochs with linearly decaying learning rate. In Figure 1 and Section 6.2, GeN is consistently performant and outperforms the baseline over multiple metrics. Somewhat surprisingly, the learning rate of GeN continues to go up even if we extend the training to 20 epochs in Figure 9, which is not captured by previous heuristic learning rate schedulers.
@@
For NLU experiments, we evaluate RoBERTa-base [38] on the GLUE [68] benchmark with LoRA, BitFit and full-parameter training (FT).
Table 4: Test performance of RoBERTa model with different methods on the GLUE benchmark (higher the better). Blue numbers are results in published papers, produced by heuristic learning rate schedulers in [28] (linear warm-up and linear decay). Red numbers are results of GeN optimizers. We report the overall (matched and mismatched) accuracy for MNLI, Matthew’s correlation for CoLA, F1 score for MRPC and QQP, and accuracy for other tasks.
@@ 6.3 Object detection & Instance Segmentation
We train a Mask R-CNN [23] with SGD on the Penn-Fudan dataset [69], following the official Pytorch tutorial which uses a linearly warm-up then constant learning rate. The loss $L(\bm{w})$ is the sum of 5 losses from the classification and the localization in different model components. The model is built on ResNet50 and pre-trained on the COCO dataset [36]. We measure the precision in Section 6.3, based on the Intersection over Union (IoU).
@@ 7 Discussion
In this work, we propose the GeN optimizers as a generally applicable and efficient approach towards the second-order optimization that is automatic and performant. Specifically, GeN is applicable to any future pre-conditioner that improves the convergence. However, additional work is needed to further our understanding of GeN. We remark that the learning rate in GeN is locally optimal but less is known about the global convergence speed. Looking forward, we expect the optimization algorithm to combine the second-order pre-conditioner (beyond AdaHessian and Sophia) and the second-order learning rate (i.e. GeN).
@@ A.3 Equivalence between finite difference and polynomial fitting
$L_{2+}$ $:=L_{0}+2\eta\mathbf{G}^{\top}\mathbf{g}+2\eta^{2}\mathbf{g}^{\top}\mathbf{H}\mathbf{g}$
@@
Then $\mathbf{G}^{\top}\mathbf{g}=\frac{L_{2+}-4L_{+}+3L_{0}}{-2\eta},\mathbf{g}^{\top}\mathbf{H}\mathbf{g}=\frac{L_{2+}-2L_{+}-L_{0}}{\eta^{2}}$, which is the equivalent to applying the second-order forward difference to $\mathbf{G}^{\top}\mathbf{g}$ and the first-order forward difference to $\mathbf{g}^{\top}\mathbf{H}\mathbf{g}$.
Hence, using the non-symmetric learning rates $(0,\eta,2\eta)$, the optimal learning rate is estimated with $O(\eta)$ error. ∎
@@ B.1 Synthetic data
In Section 5, we carefully select the best learning rate for each optimizer, through a exponential grid search from $\{10^{-k},2*10^{-k},5*10^{-k}\}$ for $k=-5,...,0$. For each optimizer, we train with each learning rate for 1000 iterations and choose the one with smallest objective value.
@@ B.2 Image Classification
For image classification problems, we use models that are pre-trained on ImageNet and can be loaded from torchvision and timm libraries. We resize all images to 224x224 and normalize the pixel values to [-1,1]. In what follows, all ResNets are trained with SGD with 0.9 momentum and 5e-4 weight decay, and all transformers (ViT, RoBERTa, GPT2) are trained with AdamW using default hyperparameters in Pytorch.
@@ B.3.1 NLG
In Figure 1, Figure 2, Figure 9 and Section 6.2, we follow the codebase of [28] and use $B=256$, sequence length 128, $\eta_{0}=1e^{-3}$, and 5 epochs. While applying, we set $\Phi=4$.
@@ B.3.2 NLU
The results of full-parameter training and BitFit training are from Table 2 of [72] and those of LoRA are from Table 2 of [28]. However, since LoRA didn’t report the F1 score of MRPC and QQP, we run LoRA under our settings (i.e., the same batch size and number of epochs) for a fair comparison. The table below records our hyper-parameters for training with GeN.
@@
While applying GeN, we set $\Phi=8$ for all tasks. For hyperparameters that were not mentioned in Table 6, we followed Table 9 of [28].
@@ B.4 Image generation
We apply GeN-Adam to pre-train a deep convolutional generative adversarial network (DCGAN) [54], following the official Pytorch tutorial, which uses a constant learning rate, with a batch size 128 and training for 5 epochs. The training uses CelebA [39], a real dataset of $>200000$ face images.
Figure 11: Left: learning of Adam over the iterations. Right: losses of the components in DCGAN.
@@ B.5 Recommendation system
We apply GeN-Adam to BERT [14] model for recommendation system, following the setting in [64], with a batch size 1024 and training for 100 epochs. We train on MovieLens-1m dataset, containing 1 million user-rating pairs. The performance is measured by the recall and Normalized Discounted Cumulative Gain (NDCG, related to the precision). The baseline is a constant learning rate given by the codebase. We observe that GeN outperforms the state-of-the-art setting (batch size 128) but needs a larger batch size to converge stably.
@@ C.1 Two modes of forward pass
As we discussed in Section 3.3, there are two modes of forward pass: with or without the activation. If the forward pass is followed by the back-propagation, then it must store the activation in each layer. The activation is $\mathbb{R}^{BTd}$ where $B$ is the batch size (say 256), $T$ is the sequence length (say 2048 for LLAMA) or pixels (say 224$\times$224 for ImageNet), and $d$ is the input dimension of one layer (say 768 for a transformer block). In practice, the activation is very expensive to store and could take up 95% of the total memory cost (see [30] Figure 3).
@@ C.2 Choice of finite difference
We recommend to use the symmetric sequences like $\{-2\xi,-\xi,0,\xi,2\xi\}$ for better precision (c.f. Section A.3). Nevertheless, this introduces a hyperparameter $\xi$ and much churn to tune it, as is the case in zero-th order optimization [43]. Therefore we use the symmetric and auto-regressive sequences that employ the previous learning rates: $(-2\eta_{t-1},-\eta_{t-1},0,\eta_{t-1},2\eta_{t-1})$. Lastly, to reduce the computation overhead, we use the fact that any quadratic function is uniquely determined by the three points. Therefore, we can use $(-\eta_{t-1},0,\eta_{t-1})$ as the shortest sequence that is both symmetric and auto-regressive.

### v3

@@ 3.4 Error analysis for optimal learning rate
We now analyze the estimation error of $\eta^{*}$ in (3.4), with batch size $B$ and mini-batch loss $\bar{L}(\bm{w}_{t})$. The estimation error consists of the sub-sampling error by using the mini-batch, and the precision error from fitting (2.3). We note that the analysis is based on the closed form of $\eta^{*}$, which is available to us through $\eta_{\text{LQA}}^{*}$ in (3.5).
@@ 4 Efficiency and Scalability Analysis
In this section, we show that GeN optimizers are as computationally efficient and scalable (in terms of the utility and the system design) as their base optimizers. We train CIFAR10 (Krizhevsky et al., 2009) on ResNet 18, 34, 50, 152 (He et al., 2016) and ViT tiny, small, base and large (Dosovitskiy et al., 2020). For fine tuning, we use the pretrained models from the PyTorch Image Models framework (Wightman, 2019).
@@
To set the stage, GeN optimizers have the same peak memory cost as the base optimizers, as all optimizers need forward passes and we adopt CPU off-loading. Hence it suffices to count the additional time complexity introduced by GeN. In the default setting, we can write
@@
where ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}F}$ is the time complexity of the forward pass, in terms of float-point operations, ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}B}$ is that of the back-propagation, and $C$ is other costs including the communication and the data loading, which are minimal on single GPU. Given that in full-parameter training, ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}B}\approx 2{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}F}$ [^5] Forward pass takes 1 unit of time. Back-propagation consists of two sub-processes – output gradient and parameter gradient, each taking 1 unit of time. See the complexity analysis in Table 1 of Bu et al. (2022)., the GeN optimizers are roughly 60% as fast as the base optimizers.
@@ 4.1 Lazy learning rate update
One simple trick to make GeN almost as fast as base optimizers is to update the learning rate every $\Phi>1$ iterations, so the additional computation is amortized. For instance, GeN achieves $86\%$ relative speed at $\Phi=4$ and $>92\%$ relative speed at $\Phi\geq 8$.
@@
In the PET setting, most parameters are frozen whose gradients are not computed. Hence ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}B}\approx{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}F}$ and the relative speed of GeN becomes $\frac{1}{1+1/\Phi}$, e.g. 50% at $\Phi=1$ and 89% at $\Phi=8$. Empirically, applying $\Phi$ has insignificant effect on the convergence.
@@ 4.2 Communication in distributed learning
For large-scale optimization, the distributed learning with multiple devices is necessary but challenging, in that (1) the communication orchestra is complicated, and (2) the communication cost is significant ($C\gg 0$). These two challenges determine if and how well an optimizer scales under a distributed solution. For data-parallel solutions such as Distributed Data Parallel and ZERO1/2 (Rajbhandari et al., 2020), the communication orchestra is relatively simple so that many Hessian-related optimizers are executable. For model-parallel and pipeline-parallel solutions such as ZERO3 and FSDP, each forward pass requires the communication of model parameters, and GeN indeed adds a noticeable volume of communication. Nevertheless, applying the lazy learning rate can essentially reduce the communication overhead to a negligible level.
@@ 6 Experiments on real data
In this section, we test GeN on real datasets across image classification, text classification, natural language generation, object detection and instance segmentation. We additionally experiment on image generation and recommendation system in Appendix B, where all training details can be found. Our experiments cover end-to-end training as well as fine-tuning, and include PET methods such as LoRA (Hu et al., ) and BitFit (Zaken et al., 2022).
@@ 6.1 Image classification
reference Krizhevsky et al. (2009) Bossard et al. (2014) Houben et al. (2013) Netzer et al. (2011) Zhou et al. (2014) ina (2021)
@@ 6.2 Natural Language Processing
For NLG experiments, we train GPT2 (Radford et al., 2019) model with LoRA on the E2E dataset (Novikova et al., 2017). We measure the quality of the generated texts by the perplexity (the exponential of loss), BLEU (Papineni et al., 2002), ROUGE-L (Lin, 2004), METEOR (Banerjee & Lavie, 2005), NIST (Doddington, 2002), and CIDEr (Vedantam et al., 2015). We use (Hu et al., ) as the baseline, which trains for 5 epochs with linearly decaying learning rate. In Figure 1 and Section 6.2, GeN is consistently performant and outperforms the baseline over multiple metrics. Somewhat surprisingly, the learning rate of GeN continues to go up even if we extend the training to 20 epochs in Figure 9, which is not captured by previous heuristic learning rate schedulers.
@@
For NLU experiments, we evaluate RoBERTa-base (Liu et al., 2019) on the GLUE (Wang et al., 2019) benchmark with LoRA, BitFit and full-parameter training (FT).
Table 4: Test performance of RoBERTa model with different methods on the GLUE benchmark (higher the better). Blue numbers are results in published papers, produced by heuristic learning rate schedulers in Hu et al. (linear warm-up and linear decay). Red numbers are results of GeN optimizers. We report the overall (matched and mismatched) accuracy for MNLI, Matthew’s correlation for CoLA, F1 score for MRPC and QQP, and accuracy for other tasks.
@@ 6.3 Object detection & Instance Segmentation
We train a Mask R-CNN (He et al., 2017) with SGD on the Penn-Fudan dataset (Wang et al., 2007), following the official Pytorch tutorial which uses a linearly warm-up then constant learning rate. The loss $L(\bm{w})$ is the sum of 5 losses from the classification and the localization in different model components. To be specific, the losses are classifier loss, bounding box regression loss in object detection, mask loss, objectness loss, and Region Proposal Network (RPN) bounding box regression loss. The model is built on ResNet50 and pre-trained on the COCO dataset (Lin et al., 2014). We measure the precision in Section 6.3, based on the Intersection over Union (IoU).
@@ 7 Discussion
In this work, we propose the GeN optimizers as a generally applicable and efficient approach towards the second-order optimization that is automatic and performant. Specifically, GeN is applicable to any future pre-conditioner that improves the convergence. However, additional work is needed to further our understanding of GeN. We remark that the learning rate in GeN is locally optimal but less is known about the global convergence speed. Looking forward, we expect further combination of more optimization algorithms and GeN, as well as exploring GeN on various domains and problems.
@@ A.3 Equivalence between finite difference and polynomial fitting
$L_{+2}$ $:=L_{0}+2\eta\mathbf{G}^{\top}\mathbf{g}+2\eta^{2}\mathbf{g}^{\top}\mathbf{H}\mathbf{g}$
@@
Then $\mathbf{G}^{\top}\mathbf{g}=\frac{L_{+2}-4L_{+}+3L_{0}}{-2\eta},\mathbf{g}^{\top}\mathbf{H}\mathbf{g}=\frac{L_{+2}-2L_{+}+L_{0}}{\eta^{2}}$, which is the equivalent to applying the second-order forward difference to $\mathbf{G}^{\top}\mathbf{g}$ and the first-order forward difference to $\mathbf{g}^{\top}\mathbf{H}\mathbf{g}$.
Hence, using the non-symmetric learning rates $(0,\eta,2\eta)$, the optimal learning rate is estimated as $\frac{\eta}{2}\frac{4L_{+}-L_{+2}-3L_{0}}{L_{+2}-2L_{+}+L_{0}}$ with $O(\eta)$ error. ∎
@@ A.4 Estimation error of $\eta^{*}$
A.5 Derivation of $\eta^{*}$
We give the derivation of (3.4), given $\{-\eta_{t-1},0,\eta_{t-1}\}$ and $\{L_{-},L_{0},L_{+}\}$. We aim to minimize (2.3): ignoring the superscript in $\mathbf{g}_{t}^{\text{optim}}$,
$\min_{\eta}L(\bm{w}_{t}-\eta\mathbf{g}_{t})\text{ or equivalently }\min_{\eta}L(\bm{w}_{t}-\eta\mathbf{g}_{t})-L(\bm{w}_{t}).$
By second-order Taylor expansion, we get
$\min_{\eta}L(\bm{w}_{t}-\eta\mathbf{g}_{t})-L(\bm{w}_{t})\approx\min_{\eta}\frac{\eta^{2}}{2}\mathbf{g}_{t}^{\top}\mathbf{H}_{t}\mathbf{g}_{t}-\eta\mathbf{G}_{t}^{\top}\mathbf{g}_{t}$
which requires the knowledge of two coefficients $\mathbf{g}_{t}^{\top}\mathbf{H}_{t}\mathbf{g}_{t}$ and $\mathbf{G}_{t}^{\top}\mathbf{g}_{t}$.
Using curve fitting, we estimate the coefficients $\mathbf{g}_{t}^{\top}\mathbf{H}_{t}\mathbf{g}_{t},\mathbf{G}_{t}^{\top}\mathbf{g}_{t}\approx A^{*},b^{*}$ by
$A^{*},b^{*}=\underset{A,b}{\text{argmin}}\sum_{\eta\in\{-\eta_{t-1},0,\eta_{t-1}\}}\left|(L(\bm{w}_{t}-\eta\mathbf{g}_{t})-L(\bm{w}_{t}))-(A\frac{\eta^{2}}{2}-b\eta)\right|^{2}$ (A.1)
The same result can be derived with finite difference as in (3.4) (though not through an optimization problem), through the numerical analysis as demonstrated in Section A.3.
Finally, we can write $\eta^{*}=b^{*}/A^{*}\approx\mathbf{G}_{t}^{\top}\mathbf{g}_{t}/\mathbf{g}_{t}^{\top}\mathbf{H}_{t}\mathbf{g}_{t}$, where the error is controlled to small values by Proposition 3.6.
As an important extension, we can use more points (say $\{-2\eta_{t-1},-\eta_{t-1},0,\eta_{t-1},2\eta_{t-1}\}$) to derive $\eta^{*}$. This usually gives a slightly more stable convergence, almost the same accuracy, but more computational overhead.
Using curve fitting, we easily extend to
$\mathbf{g}_{t}^{\top}\mathbf{H}_{t}\mathbf{g}_{t},\mathbf{G}_{t}^{\top}\mathbf{g}_{t}\approx\underset{A,b}{\text{argmin}}\sum_{\eta\in\{-2\eta_{t-1},-\eta_{t-1},0,\eta_{t-1},2\eta_{t-1}\}}\left|(L(\bm{w}_{t}-\eta\mathbf{g}_{t})-L(\bm{w}_{t}))-(A\frac{\eta^{2}}{2}-b\eta)\right|^{2}$
Using finite difference, the derivation is more complicated than (3.4),
$\eta^{*}\approx\eta\cdot\frac{-\frac{1}{12}L_{+2}+\frac{2}{3}L_{+}-\frac{2}{3}L_{-}+\frac{1}{12}L_{-2}}{-\frac{1}{12}L_{+2}+\frac{4}{3}L_{+}-\frac{5}{2}L_{0}+\frac{4}{3}L_{-}-\frac{1}{12}L_{-2}}$ (A.2)
where $L_{\pm 2}=L(\bm{w}_{t}\pm 2\eta_{t-1}\mathbf{g}_{t})$. We refer to https://en.wikipedia.org/wiki/Finite_difference_coefficient for the table of finite difference coefficients.
@@ B.1 Synthetic data
In Section 5, we carefully select the best learning rate for each optimizer, through a exponential grid search from $\{10^{-k},2*10^{-k},5*10^{-k}\}$ for $k=5,4,...,0$. For each optimizer, we train with each learning rate for 1000 iterations and choose the one with smallest objective value.
@@ B.2 Image Classification
For image classification problems, we use models that are pre-trained on ImageNet and can be loaded from torchvision and timm libraries. We resize all images to 224x224 and normalize the pixel values to [-1,1]. In what follows, all ResNets are trained with SGD with 0.9 momentum and 5e-4 weight decay, and ViT is trained with AdamW using default hyperparameters in Pytorch.
@@ B.3 Natural Language Processing
All transformers (RoBERTa, GPT2) are trained with AdamW using default hyperparameters in Pytorch.
@@ B.3.1 NLG
In Figure 1, Figure 2, Figure 9 and Section 6.2, we follow the codebase of Hu et al. and use $B=256$, sequence length 128, $\eta_{0}=1e^{-3}$, and 5 epochs. While applying, we set $\Phi=4$.
@@ B.3.2 NLU
The results of full-parameter training and BitFit training are from Table 2 of Zaken et al. (2022) and those of LoRA are from Table 2 of Hu et al. . However, since LoRA didn’t report the F1 score of MRPC and QQP, we run LoRA under our settings (i.e., the same batch size and number of epochs) for a fair comparison. The table below records our hyper-parameters for training with GeN.
@@
While applying GeN, we set $\Phi=8$ for all tasks. For hyperparameters that were not mentioned in Table 6, we followed Table 9 of Hu et al. .
@@ B.4 Image generation
We apply GeN-Adam to pre-train a deep convolutional generative adversarial network (DCGAN) Radford et al. (2015), following the official Pytorch tutorial, which uses a constant learning rate, with a batch size 128 and training for 5 epochs. The training uses CelebA Liu et al. (2015), a real dataset of $>200000$ face images.
Figure 11: Left: learning rate of GeN-Adam over the iterations. Right: losses of the components in DCGAN, where D is the discriminator and G is the generator.
@@ B.5 Recommendation system
We apply GeN-Adam to BERT Kenton & Toutanova (2019) model for recommendation system, following the setting in Sun et al. (2019), with a batch size 1024 and training for 100 epochs. We train on MovieLens-1m dataset, containing 1 million user-rating pairs. The performance is measured by the recall and Normalized Discounted Cumulative Gain (NDCG, related to the precision). The baseline is a constant learning rate given by the codebase. We observe that GeN outperforms the state-of-the-art setting (batch size 128) but needs a larger batch size to converge stably.
@@ C.1 Two modes of forward pass
As we discussed in Section 3.3, there are two modes of forward pass: with or without the activation. If the forward pass is followed by the back-propagation, then it must store the activation in each layer. The activation is $\mathbb{R}^{BTd}$ where $B$ is the batch size (say 256), $T$ is the sequence length (say 2048 for LLAMA) or pixels (say 224$\times$224 for ImageNet), and $d$ is the input dimension of one layer (say 768 for a transformer block). In practice, the activation is very expensive to store and could take up 95% of the total memory cost (see Jain et al. (2020) Figure 3).
@@ C.2 Choice of finite difference
We recommend to use the symmetric sequences like $\{-2\xi,-\xi,0,\xi,2\xi\}$ for better precision (c.f. Section A.3). Nevertheless, this introduces a hyperparameter $\xi$ and much churn to tune it, as is the case in zero-th order optimization Malladi et al. (2023). Therefore we use the symmetric and auto-regressive sequences that employ the previous learning rates: $(-2\eta_{t-1},-\eta_{t-1},0,\eta_{t-1},2\eta_{t-1})$. Lastly, to reduce the computation overhead, we use the fact that any quadratic function is uniquely determined by the three points. Therefore, we can use $(-\eta_{t-1},0,\eta_{t-1})$ as the shortest sequence that is both symmetric and auto-regressive.